OpenStudy (anonymous):

correct or wrong 1 (-6) +1 (-3) -1=8?

OpenStudy (amistre64):

1(-6) = -6 1(-3) = -3 -1 = -1 ---------- =-10

OpenStudy (anonymous):

it's wrong... -6-3-1=-10 not 8....

OpenStudy (anonymous):

-10 is the answer

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ic ic 6 question to go i submit my test =) haha thx people